OpenStudy (anonymous):

How would I solve this? I know what the answer is but I can't seem to get the steps right... Can someone give me a brief overview on how to do it?(x+(3+5i))^2

hartnn (hartnn):

\(\Large x+3+5i \) Real part = A = x+3 Imaginary part = B = 5 \(\Large (A+iB)^2 = A^2 +2iB -B^2\) try it ? :)

OpenStudy (anonymous):

Can you explain how you got that?

OpenStudy (anonymous):

Can someone help please?